Click here to read the eBook: Business and Financial Risk
Click here to read the eBook: Determining the Optimal Capital Structure

BREAKEVEN AND LEVERAGE

Wingler Communications Corporation (WCC) produces premium stereo headphones that sell for $28.70 per set, and this year's sales are expected to be 450,000 units. Variable production costs for the expected sales under present production methods are estimated at $10,500,000, and fixed production (operating) costs at present are $1,560,000. WCC has $4,800,000 of debt outstanding at an interest rate of 8%. There are 240,000 shares of common stock outstanding, and there is no preferred stock. The dividend payout ratio is 70%, and WCC is in the 40% federal-plus-state tax bracket.

The company is considering investing $7,200,000 in new equipment. Sales would not increase, but variable costs per unit would decline by 20%. Also, fixed operating costs would increase from $1,560,000 to $1,800,000. WCC could raise the required capital by borrowing $7,200,000 at 10% or by selling 240,000 additional shares of common stock at $30 per share.

  1. What would be WCC's EPS (1) under the old production process, (2) under the new process if it uses debt, and (3) under the new process if it uses common stock? Do not round intermediate calculations. Round your answers to the nearest cent.
    1.  $
    2.  $
    3.  $

  2. At what unit sales level would WCC have the same EPS assuming it undertakes the investment and finances it with debt or with stock? {Hint: V = variable cost per unit = $8,400,000/450,000, and EPS = [(PQ - VQ - F - I)(1 - T)]/N. Set EPSStock = EPSDebt and solve for Q.} Do not round intermediate calculations. Round your answer to the nearest whole _____units

  3. At what unit sales level would EPS = 0 under the three production/financing setups - that is, under the old plan, the new plan with debt financing, and the new plan with stock financing? (Hint: Note that VOld = $10,500,000/450,000, and use the hints for part b, setting the EPS equation equal to zero.) Do not round intermediate calculations. Round your answers to the nearest whole.
    Old plan ____ units
    New plan with debt financing ____ units
    New plan with stock financing _____ units

  4. On the basis of the analysis in parts a through c, and given that operating leverage is lower under the new setup, which plan is the riskiest, which has the highest expected EPS, and which would you recommend? Assume here that there is a fairly high probability of sales falling as low as 250,000 units, and determine EPSDebt and EPSStock at that sales level to help assess the riskiness of the two financing plans. Do not round intermediate calculations. Round your answers to two decimal places. Negative amount should be indicated by a minus sign.
    EPSDebt = $_______
    EPSStock = $_______

    The input in the box below will not be graded, but may be reviewed and considered by your instructo
Answer #1

Слоу veriable cost Ausha e 2. » ч гъъъъ . (29) SUVъ с тобо0 0 0 contributies - 2415000 fixed cost = (1560000) 8 сбо 60 Г 8656

6 т е 221 СЫ е д) Ә2 объъ ко) 2016 О са 5+ - - 38yo 1 Т 16) БЪы е- чээ1 г 6 учу оо енг а 666 oы ner. of shares 29 653 eß Ч-01

Sale bolo [ T Aus @ els in stock opties & Els in Debt Price xo)-VC - FC = 1) (St.xa)-rc-F-JC 1-7) No. of Shares ao at Shay (2

ane a otel plany ( Sa - VR -F - I )(- D o (28.75 -2.330 - 1560000 1-0.40) 24000 126420 – 1560000) x 0.60 ao 26:42 a = 1560000

Ans e a geen plan of saper (98.45Q -1.866 -1 800000 - 3846096 48oooo 26.89 0 – 2184000 - s. i o = 81219.78 vuit b- Aus g Post2 12 Stъо (2262) And a fast less in stock eption. Se 2 Sъбvo x 28. до C V., 2 Jхоо х 2-77 contributien - Fixed est орт Т » Mt